# Computational Complexity of Sufficiency in Decision Problems

### v2

@@ 9.1 Computational Decision Theory
The complexity of decision-making has been studied extensively. Papadimitriou [10] established foundational results on the complexity of game-theoretic solution concepts. Our work extends this to the meta-question of identifying relevant information. For a modern treatment of complexity classes, see Arora and Barak [3].
Closest prior work and novelty.
Closest to our contribution is the literature on feature selection and model selection hardness, which proves NP-hardness of selecting informative feature subsets and inapproximability for minimum feature sets [4, 2]. Those results analyze predictive relevance or compression objectives. We study decision relevance and show coNP-completeness for sufficiency checking, a different quantifier structure with distinct proof techniques and a full hardness/tractability landscape under explicit encoding assumptions, mechanized in Lean 4. The formalization aspect is also novel: prior work establishes hardness on paper, while we provide machine-checked reductions with explicit polynomial bounds.
@@ 9.2 Feature Selection
In machine learning, feature selection asks which input features are relevant for prediction. This is known to be NP-hard in general [4]. Our results show the decision-theoretic analog is coNP-complete for both checking and minimization.
@@ 9.3 Value of Information
The value of information (VOI) framework [7] quantifies the maximum rational payment for information. Our work addresses a different question: not the value of information, but the complexity of identifying which information has value.
@@ 9.4 Model Selection
Statistical model selection (AIC [1], BIC [13], cross-validation [15]) provides practical heuristics for choosing among models. Our results provide theoretical justification: optimal model selection is intractable, so heuristics are necessary.
@@ 10 Conclusion
Role of LLMs in this work. This paper was developed through human-AI collaboration. The author provided the core intuitions—the connection between decision-relevance and computational complexity, the conjecture that SUFFICIENCY-CHECK is coNP-complete, and the insight that the $\Sigma_{2}^{\textsf{P}}$ structure collapses for MINIMUM-SUFFICIENT-SET. Large language models (Claude, GPT-4) served as implementation partners for proof drafting, Lean formalization, and generation.
@@
The proofs are machine-checked; their validity is independent of generation method. We disclose this methodology in the interest of academic transparency.
@@
The reduction constructions and key equivalence theorems are machine-checked in Lean 4 (see Appendix A for proof listings). The formalization verifies that the TAUTOLOGY reduction correctly maps tautologies to sufficient coordinate sets; complexity classifications (coNP-completeness, $\Sigma_{2}^{\textsf{P}}$-completeness) follow by composition with standard complexity-theoretic results (TAUTOLOGY is coNP-complete, $\exists\forall$-SAT is $\Sigma_{2}^{\textsf{P}}$-complete). The strengthened gadget showing that non-tautologies yield instances with all coordinates relevant is also formalized.
@@
    Exact bounds. SUFFICIENCY-CHECK is coNP-complete—both coNP-hard and in coNP.
@@
    Encoding-regime separation. Under the explicit-state encoding, SUFFICIENCY-CHECK is polynomial in $|S|$. Under ETH, there exist succinctly encoded worst-case instances (witnessed by a strengthened gadget family with $k^{*}=n$) requiring $2^{\Omega(n)}$ time. Intermediate regimes are not ruled out by the lower-bound statement.
The Complexity Conservation Law
Section 8 develops a quantitative consequence: when a problem requires $k$ dimensions and a model handles only $j<k$ natively, the remaining kjk-j dimensions must be handled externally at each decision site. For $n$ sites, total external work is $(k-j)\times n$.
This conservation law is formalized in Lean 4 (HardnessDistribution.lean), proving:
  • Conservation: complexity cannot be eliminated, only redistributed
@@
Several questions remain for future work:
  • Fixed-parameter tractability: Is SUFFICIENCY-CHECK FPT when parameterized by the size of the minimal sufficient set?
@@
  • Quantum complexity: Does quantum computation provide speedups for sufficiency checking?
@@
Practical interpretation
While this work is theoretical, the complexity bounds have practical meaning. Under ETH, worst-case instances of SUFFICIENCY-CHECK and related problems grow exponentially with problem size; practitioners should therefore not expect polynomial‑time worst‑case solvers for arbitrary inputs. In practice, real datasets often have structure (sparsity, bounded dependency depth, separable utilities) that the tractable subcases capture. Thus, engineering efforts should focus on (a) detecting and exploiting structural restrictions in inputs, and (b) designing heuristic or approximate methods for large unstructured instances. Our experiments illustrate performance on representative inputs, but the theoretical bounds underline the need for domain constraints to achieve scalable, worst‑case guarantees.

### v3

@@ 10 Related Work
10.1 Formalized Complexity and Mechanized Proof Infrastructure
Formalizing complexity-theoretic arguments in proof assistants remains comparatively sparse. Lean and Mathlib provide the current host environment for this artifact [16, 17, 67]. Closest mechanized precedents include verified computability/complexity developments in Coq and Isabelle [21, 32, 45, 44, 26], and certifying toolchain work that treats proofs as portable machine-checkable evidence [34, 43].
Recent LLM-evaluation work also highlights inference nondeterminism from numerical precision and hardware configuration, reinforcing the value of proof-level claims whose validity does not depend on stochastic inference trajectories [72].
@@ 10.2 Computational Decision Theory
The complexity of decision-making has been studied extensively. Foundational treatments of computational complexity and strategic decision settings establish the baseline used here [47, 5]. Our work extends this to the meta-question of identifying relevant information. Decision-theoretic and information-selection framing in this paper also sits in the tradition of statistical sufficiency and signaling/information economics [20, 65, 42, 40, 29].
Large-sample Bayesian-network structure learning and causal identification are already known to be hard in adjacent formulations [13, 61, 30]. Our object differs: we classify coordinate sufficiency for optimal-action invariance, with mechanized reductions and regime typing.
Relation to prior hardness results.
Closest adjacent results are feature-selection/model-selection hardness results for predictive subset selection [9, 3]. Relative to those works, this paper changes two formal objects: (i) the reductions are machine-checked (TAUTOLOGY and $\exists\forall$-SAT mappings with explicit polynomial bounds), and (ii) the output is a regime-typed hardness/tractability map for decision relevance under explicit encoding assumptions. The target predicate here is decision relevance, not predictive compression.
10.3 Succinct Representations and Regime Separation
Representation-sensitive complexity is established in planning and decision-process theory: classical and compactly represented MDP/planning problems exhibit sharp complexity shifts under different input models [48, 41, 37]. The explicit-vs-succinct separation in this paper is the same representation-sensitive phenomenon for the coordinate-sufficiency predicate.
The distinction in this paper is the object and scope of the classification: we classify decision relevance (sufficiency, anchor sufficiency, and minimum sufficient sets) for a fixed decision relation, with theorem-level regime typing and mechanized reduction chains.
10.4 Oracle and Query-Access Lower Bounds
Query-access lower bounds are a standard source of computational hardness transfer [18, 11]. Our [Q_fin] and [Q_bool] results are in this tradition, but specialized to the same sufficiency predicate used throughout the paper: they establish access-obstruction lower bounds while keeping the underlying decision relation fixed across regimes.
[…]
10.5 Energy Landscapes and Chemical Reaction Networks
The geometric interpretation of decision complexity connects to the physics of energy landscapes. Wales established a widely used picture in which molecular configurations form high-dimensional landscapes with basins and transition saddles [68, 69, 70]. Protein-folding theory uses the same landscape language via funnel structure toward native states [46]. In this paper’s formal model, the corresponding core object is coordinate-projection invariance of the optimizer map: sufficiency asks whether fixing coordinates preserves Opt, while insufficiency is witnessed by agreeing projections with different optimal sets (Definition 2.4, Proposition 2.6).[C:D2.4;P2.6;R:DM](L: DP7,DP8.)
Section 2.7 makes this transfer rule theorem-level: encoded physical claims are derived from core claims through an explicit assumption-transfer map, and an encoded physical counterexample refutes the corresponding core rule on that encoded slice.[C:P2.34;C2.35;R:AR](L: CT6,CT5,CT8,CT7.)
Under that mapping, the explicit-vs-succinct split is the mechanized statement of access asymmetry: [E] admits polynomial verification on explicit state access, while [S+ETH] admits worst-case exponential lower bounds on succinct inputs (Theorem 5.1); finite query access already exhibits indistinguishability barriers (Proposition 5.3). The complexity separation is mechanized.[C:T5.1;P5.3;R:E,S+ETH,Q_fin](L: CC16,CC23,CC24,CC50.)
The geometric core is mechanized directly: product-space slice cardinality, hypercube node-count identity, and node-vs-edge separation are proved in Lean at the decision abstraction level, with edge witnesses tied exactly to non-sufficiency.(L: DQ3, L25, L68, DP21.)
The circuit-to-decision instantiation is also mechanized in a separate bridge layer: geometry and dynamics are represented as typed components of a circuit object; decision circuits add an explicit interpretation layer; molecule-level constructors instantiate both views.(L: DQ519, DQ518, DQ517, DQ155, DQ154, DQ159, DQ160.)
Chemical reaction networks provide a concrete physical class where this encoding applies directly. Prior work establishes hardness for output optimization, reconfiguration, and reconstruction in CRN settings [4, 2, 19]. Those domain-specific CRN complexity results are external literature claims, not mechanized in this Lean artifact. The mechanized statement here is the transfer principle: once a CRN decision task is encoded as a C1–C4 decision problem, it inherits the same sufficiency hardness/tractability regime map (Theorems 7.2, 4.6, 6.1).[C:T7.2,4.6,6.1;R:E,S,TA](L: CR1,CC61,CC69,CC70,CC71,CC72.)
The physical bridge is not a single-domain add-on: Theorem 2.37 composes law-objective semantics, reduction equivalence, and hardness lower-bound accounting in one mechanized theorem, so physical instantiation and complexity claims share one proof chain.[C:T2.37;R:AR](L: CT4.)
For matter-only dynamics, the law-induced objective specialization is mechanized for arbitrary universes (arbitrary state/action types and transition-feasibility relations): the utility schema is induced from transition laws, and under a strict allowed-vs-blocked gap with nonempty feasible set, the optimizer equals the feasible-action set; with logical determinism at the action layer, the optimizer is singleton.[C:P2.36;R:AR](L: DQ8, DQ5, DQ7, DQ6.)
The interface-time semantics are also explicit and mechanized: timed states are $\mathbb{N}$-indexed, one decision event is exactly one tick, run length equals elapsed interface time, and this unit-time law is invariant across substrate tags under interface-consistency assumptions.[C:P2.28,2.29,2.30,2.31;R:AR](L: DQ548,DQ537,DQ539,DQ534,DQ541,DQ549.)
Regime-side interface machinery is also formalized as typed access objects with an explicit simulation preorder and certificate-upgrade statements. The access-channel and five-way composition theorems are encoded as assumption-explicit composition laws rather than informal metaphors.[C:P2.34;T2.37;R:AR](L: PC10, PC12, PC15, PC17, PC19, PC20, PC21.)
10.6 Thermodynamic Costs and Landauer’s Principle
The thermodynamic lift in Section 5 converts declared bit-operation lower bounds into declared energy/carbon lower bounds via linear conversion constants. In this artifact, those conversion consequences are mechanized conditionally through Propositions 5.28, 5.29, 5.30, and 5.31. Landauer’s principle and reversible-computation context [33, 7] define the physical assumption set; the constants are declared premises, not Lean-derived constants.[C:P5.28,5.29,5.30,5.31;R:AR](L: CC65,CC66,CC67,CC68,CC64.)
Coupling this to integrity yields a mechanized policy condition: when exact competence is hardness-blocked in the declared regime, integrity requires abstention or explicitly weakened guarantees (Propositions 3.61, 3.60). A stronger behavioral claim such as “abstention is globally utility-optimal for every objective” is not currently mechanized and remains objective-dependent.[C:P3.61,3.60;R:AR](L: IC26,IC24,CC30,IC27,IC7,IC6.)
@@ 10.7 Feature Selection
In machine learning, feature selection asks which input features are relevant for prediction. This is known to be NP-hard in general [9, 25]. Our results show the decision-theoretic analog is coNP-complete for both checking and minimization. [C:T4.6,4.7;R:S](L: CC61,CC36.)
@@ 10.8 Value of Information
The value of information (VOI) framework [27, 52] quantifies the maximum rational payment for information. Our work addresses a different question: not the value of information, but the complexity of identifying which information has value.
@@ 10.9 Model Selection
Statistical model selection (AIC/BIC/cross-validation) provides practical heuristics for choosing among models [1, 56, 66]. Our results formalize the regime-level reason heuristic selection appears: without added structural assumptions, exact optimal model selection inherits worst-case intractability, so heuristic methods implement explicit weakened-guarantee policies for unresolved structure. [C:T4.6,5.1,6.1;R:E,S,S+ETH](L: CC61,CC16,CC69,CC70,CC71.)
10.10 Information Compression vs Certification Length
The paper now makes an explicit two-part information accounting object (Definition 3.45): raw report bits and evidence-backed certification bits. This sharpens the information-theoretic reading of typed claims: compression of report payload is not equivalent to compression of certifiable truth conditions [57, 14, 64].
From a zero-error viewpoint, this distinction is structural rather than stylistic: compressed representations can remain short while exact decodability depends on confusability structure and admissible decoding conditions [58, 31, 38, 15]. The same separation appears here between short reports and evidence-backed exact claims.
The rate-distortion and MDL lines make the same split in different language: there is a difference between achievable compression rate, computational procedure for obtaining/optimizing that rate, and certifiable model adequacy [59, 8, 54, 24, 35].
Formally, Theorem 3.47 and Corollary 3.49 characterize the exact-report boundary: admissible exact reports require a strict certified-bit gap above raw payload, while hardness-blocked exact reports collapse to raw-only accounting. This directly links representational succinctness to certification cost under the declared contract, rather than treating “short reports” as automatically high-information claims. [C:T3.47;P3.46;C3.49;R:AR](L: CC18,CC11,CC21.)
10.11 Certifying Outputs and Proof-Carrying Claims
Our integrity layer matches the certifying-algorithms pattern: algorithms emit candidate outputs together with certificates that can be checked quickly, separating producing claims from verifying claims [39, 43]. In this paper, Definition 3.8 is exactly that soundness discipline.
At the systems level, this is the same architecture as proof-carrying code: a producer ships evidence and a consumer runs a small checker before accepting the claim [43, 39]. Our competence definition adds the regime-specific coverage/resource requirement that certifying soundness alone does not provide.
The feasibility qualifier in Definition 3.9 yields the bounded-rationality constraint used in this paper: admissible policy is constrained by computational feasibility under the declared resource model (Remark 2.25) [71, 5].[C:P3.58;R:AR] (L: IC18.)
Cryptographic-verifiability perspective (scope).
The role of cryptography in this paper is structural verifiability, not secrecy: the relevant analogy is certificate-carrying claims with lightweight verification, not confidential encoding [23, 43]. Concretely, the typed-report discipline is modeled as a game-based reporting interface: a prover emits $(r,\pi)$ and a verifier accepts exactly when $\pi$ certifies the declared report type under the active contract. In that model, Evidence–Admissibility Equivalence gives completeness for admissible report types, while Typed Claim Admissibility plus Exact Claims Require Exact Evidence give soundness against exact overclaiming. Certified-confidence constraints then act as evidence-gated admissibility constraints on top of that verifier interface, so raw payload bits and certified bits represent distinct resources rather than stylistic notation variants. [C:P3.18,3.16,3.21,3.23;R:AR](L: IC2,CC75,IC35,IC38.)
Three-axis integration.
In the cited literature, these pillars are treated separately: representation-sensitive hardness, query-access lower bounds, and certifying soundness disciplines. This paper composes all three for the same decision-relevance object in one regime-typed and machine-checked framework. [C:T5.1;P5.3,3.16;R:S+ETH,Qf,AR](L: CC16,CC50,CC75.)
@@ 11 Conclusion
Role of LLMs in this work. This paper was developed through human-AI collaboration. The author provided the core intuitions: the connection between decision-relevance and computational complexity, the conjecture that SUFFICIENCY-CHECK is coNP-complete, and the insight that the $\Sigma_{2}^{\textsf{P}}$ structure collapses for MINIMUM-SUFFICIENT-SET. Large language models (Claude, GPT-4) served as implementation partners for proof drafting, Lean formalization, and generation.
@@
The proofs are machine-checked; their validity is independent of generation method. This methodology is disclosed for academic transparency.
@@
  • Full intermediate query-access lower bounds are formalized as a finite-state Opt-oracle core ($\Omega(|S|)$, Boolean instantiation $\Omega(2^{n})$) plus Boolean-interface refinements for value-entry and state-batch access, with explicit subproblem-to-full transfer, weighted randomized robustness, and oracle-lattice transfer/strictness closures (Propositions 5.35.16)
@@
  • Discrete-time interface semantics are formalized: time is $\mathbb{N}$-valued, decision events are exactly one-tick transitions, run length equals elapsed time, and substrate-consistent realizations preserve the same one-unit event law (Propositions 2.282.31)
  • Physical-system transport is formalized: encoded physical claims are derived from core rules through explicit assumption transfer, physical counterexamples map to encoded core failures, and the bundled bridge theorem composes law semantics, reduction equivalence, and required-work lower bounds (Proposition 2.34, Corollary 2.35, Theorem 2.37)
[C:T4.6,4.7,4.9,5.1,6.1,2.37;P5.3,5.7,5.11,5.8,5.15,5.16,2.28,2.29,2.30,2.31,2.34;C2.35;R:E,S,S+ETH,Qf,Qb,AR] (L: CC61,CC36,CC5,CC16,CC69,CC50,L9,L8,L11,L21,L20,DQ548,DQ537,DQ539,DQ541,CT6,CT8,CT4.)
@@
Beyond classification, the paper contributes a formal claim-typing framework (Section 3): structural complexity is a property of the fixed decision question, while representational hardness is regime-conditional access cost. This is why encoding-regime changes can move practical hardness without changing the underlying semantics. [C:P2.20;T3.4;R:AR] (L: CC62,CC76.)
The reduction constructions and key equivalence theorems are machine-checked in Lean 4 (see Appendix A for proof listings). The formalization verifies that the TAUTOLOGY reduction correctly maps tautologies to sufficient coordinate sets; complexity classifications (coNP-completeness, $\Sigma_{2}^{\textsf{P}}$-completeness) follow by composition with standard complexity-theoretic results (TAUTOLOGY is coNP-complete, $\exists\forall$-SAT is $\Sigma_{2}^{\textsf{P}}$-complete). The strengthened gadget showing that non-tautologies yield instances with all coordinates relevant is also formalized. [C:T4.6,4.7,4.9,5.1;R:S,S+ETH] (L: CC61,CC36,CC5,CC16,CC24.)
@@
    Exact bounds. SUFFICIENCY-CHECK is coNP-complete, both coNP-hard and in coNP.
@@
    Encoding-regime separation. Under [E], SUFFICIENCY-CHECK is polynomial in $|S|$. Under [S+ETH], there exist succinct worst-case instances (with $k^{*}=n$) requiring $2^{\Omega(n)}$ time. Under [Q_fin], the Opt-oracle core has $\Omega(|S|)$ worst-case query complexity (Boolean instantiation $\Omega(2^{n})$), and under [Q_bool] the value-entry/state-batch refinements preserve the obstruction with weighted-cost transfer closures.
[C:T4.6,5.1;P5.3,5.7,5.11,5.8;R:E,S,S+ETH,Qf,Qb] (L: CC61,CC16,CC50,L9,L8,L11.)
The Complexity Redistribution Corollary
Section 9 develops a quantitative consequence: when a problem requires $k$ dimensions and a model handles only $j<k$ natively, the remaining kjk-j dimensions must be handled externally at each decision site. For $n$ sites, total external work is $(k-j)\times n$.[C:T9.5;R:LC] (L: HD21.)
The set identity is elementary; its operational content comes from composition with the hardness results on exact relevance minimization. This redistribution corollary is formalized in Lean 4 (HardnessDistribution.lean), proving:
  • Redistribution identity: complexity burden cannot be eliminated by omission, only moved between native handling and external handling
@@
The landscape above is complete for the static sufficiency class under C1–C4 and the declared regimes; the items below are adjacent-class extensions or secondary refinements. Several questions remain for future work:
  • Fixed-parameter tractability (primary): Is SUFFICIENCY-CHECK FPT when parameterized by the minimal sufficient-set size $k^{*}$, or is it W[2]-hard under this parameterization?
[…]
@@
  • Typed-confidence calibration: For signaled reports $(r,g,p_{\mathrm{self}},p_{\mathrm{cert}})$, characterize proper scoring and calibration objectives that preserve signal consistency and typed admissibility.
@@
Practical Corollaries
The practical corollaries are regime-indexed and theorem-indexed:
  • [E] and structured regimes: polynomial-time exact procedures exist (Theorem 6.1).
  • [Q_fin]/[Q_bool] query-access lower bounds: worst-case Opt-oracle complexity is $\Omega(|S|)$ in the finite-state core (Boolean instantiation $\Omega(2^{n})$), and value-entry/state-batch interfaces satisfy the same obstruction in the mechanized Boolean refinement (Propositions 5.35.16), with randomized weighted robustness and oracle-lattice closures.
  • [S+ETH] hard families: exact minimization inherits exponential worst-case cost (Theorem 5.1 together with Theorem 4.6).
  • [S_bool] mechanized criterion: minimization reduces to relevance-cardinality constraints (Corollary 8.6).
  • Redistribution consequences: omitted native coverage externalizes work with explicit growth/amortization laws (Theorems 9.39.6).
  • Physical-system instantiation: once a physical class is encoded into the declared decision interface, the same regime-typed hardness/tractability map and integrity constraints apply to that class (Section 2.7).[C:P2.34;C2.35;T2.37;R:AR](L: CT6,CT8,CT4.)
Hence the design choice is typed: enforce a tractable regime, or adopt weakened guarantees with explicit verification boundaries.[C:T6.1,5.1;R:RG](L: CC71,CC16.) Equivalently, under the attempted-competence matrix, over-specification is rational only in attempted-competence-failure cells; once exact competence is available in the active regime (or no attempted exact competence was made), persistent over-specification is irrational for the same verified-cost objective.[C:P3.60;R:AR](L: IC27.) By Proposition 3.60, the integrity-preserving matrix contains 3 irrational cells and 1 rational cell.[C:P3.60;R:AR](L: IC27.)
Epistemic Alignment Implications for Agent Training
Within this paper’s formalism, epistemic alignment is a claim-discipline property: outputs are asserted only when certifiable, and otherwise emitted as abstain reports (optionally carrying tentative guesses and self-reflected confidence under zero certified confidence). Formally, this is the integrity/competence split of Section 3, not a full value-alignment claim.[C:D3.8,3.9;P3.58;R:TR](L: IC18.)
Truth-Objective Abstention Requirement.
If a solver-agent declares a truth-preserving objective for claims about relation $\mathcal{R}$ in regime $\Gamma$, uncertified outputs are admissible only as $\mathsf{ABSTAIN}/\mathsf{UNKNOWN}$ rather than asserted answers. Under the signal discipline, this includes abstain-with-guess/self-confidence reports while keeping certification evidence-gated.[C:D3.19,3.20;P3.23,3.29;C3.25;R:AR](L: IC46,IC38,IC41.) Equivalently, training policies that suppress abstention in uncertified regions violate the typed claim discipline under the declared task model.[C:D3.8;P3.58,3.61,3.60;R:AR](L: IC18,CC30,IC27.)
This yields a training doctrine for solver-agents under declared tasks/regimes:
  1. 1.
    Treat $\mathsf{ABSTAIN}/\mathsf{UNKNOWN}$ as first-class admissible outputs, not errors.[C:D3.8;R:TR]
  2. 2.
    For signaled reports, maintain two channels: self-reflected confidence may be emitted, but certified confidence is positive only with evidence.[C:D3.20;P3.23,3.30;R:AR](L: IC38,IC47.)
  3. 3.
    Treat runtime disclosure as scalar-witnessed: always emit exact step-count, and emit completion percentage only when a declared bound exists.[C:P3.26,3.27,3.28;R:AR](L: IC42,IC35,IC44.)
  4. 4.
    Optimize competent coverage subject to integrity constraints, rather than optimizing answer-rate alone.[C:D3.9;P3.58;R:AR](L: IC18.)
  5. 5.
    In hardness-blocked regimes, abstention or explicit guarantee-weakening is structurally required unless regime assumptions are changed.[C:T5.1;P3.61,3.60;R:S,S+ETH](L: CC16,CC30,IC27.)
Operationally, a reporting schema consistent with this framework tracks at least five quantities: integrity violations, competent non-abstaining coverage on the declared scope, abstention quality under regime shift, the calibration gap between self-reflected and certified confidence channels, and scalar runtime witness distributions (with bound-qualified fraction semantics when bounds are declared). This is consistent with bounded-rationality feasibility discipline under declared resource models [71, 5].